Find the tensions in the ropes for each case. Note that theta 1= 35, theta 2= 55, theta 3=60, m1= 6 kg, and m2=9 kg


1
Expert's answer
2021-02-19T10:30:28-0500

The picture must be given, so I imagine the picture.



1) "\theta_1 = 35^o \space m_1 = 6kg"

"(m_1g)^2 = 2T^2 - 2T^2cos\theta_1 = 2T^2(1-cos\theta_1) \to T = \large\frac{m_1g}{\sqrt{2(1-cos\theta_1)}}" = 99.7N

2) "\theta_1 = 55^o \space m_1 = 6kg"

"(m_1g)^2 = 2T^2 - 2T^2cos\theta_1 = 2T^2(1-cos\theta_1) \to T = \large\frac{m_1g}{\sqrt{2(1-cos\theta_1)}}" = 64.9N

3) "\theta_1 = 60^o \space m_1 = 6kg"

"(m_1g)^2 = 2T^2 - 2T^2cos\theta_1 = 2T^2(1-cos\theta_1) \to T = \large\frac{m_1g}{\sqrt{2(1-cos\theta_1)}}" = 60N

4) "\theta_1 = 35^o \space m_1 = 9kg"

"(m_1g)^2 = 2T^2 - 2T^2cos\theta_1 = 2T^2(1-cos\theta_1) \to T = \large\frac{m_1g}{\sqrt{2(1-cos\theta_1)}}" = 149.6N

5) "\theta_1 = 55^o \space m_1 = 9kg"

"(m_1g)^2 = 2T^2 - 2T^2cos\theta_1 = 2T^2(1-cos\theta_1) \to T = \large\frac{m_1g}{\sqrt{2(1-cos\theta_1)}}" = 97.4N

6) "\theta_1 = 60^o \space m_1 = 9kg"

"(m_1g)^2 = 2T^2 - 2T^2cos\theta_1 = 2T^2(1-cos\theta_1) \to T = \large\frac{m_1g}{\sqrt{2(1-cos\theta_1)}}" = 90N